OpenStudy (anonymous):

For these matrices: S=[u,v,-v,u] T=[x,y,-y,x] NOTE: these matrices are 2x2 matrices, the first two on the top row and the last two on the bottom row. (sorry I can't express them properly here). a) Find the determinants of S, T and ST b) Use ordinary algebra to verify that det(S)det(T)=det(ST) c) Express (7^2 + 2^2)(4^2 + 1^2) as a sum of squares of two integers.

OpenStudy (dumbcow):

for any 2X2 matrix a b c d det = ad - bc

myininaya (myininaya):

give me just a sec k?

myininaya (myininaya):

i will scan what i have so far... it isn't complete

OpenStudy (anonymous):

thanks everyone - especially myininaya :)

myininaya (myininaya):

i was fixing to type this i think i will just scan it lol

OpenStudy (anonymous):

good idea, much easier :)

myininaya (myininaya):

you know a^2+2ab+b^2=(a+b)^2 and a^2-2ab+b^2=(a-b)^2 right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

do you see i add in a zero so I could write (ux-vy)^2+(vx+yu)^2

OpenStudy (anonymous):

yes indeed :) thank you so much!!!

myininaya (myininaya):

k cool are there any questions? please let me know if you don't understand something

OpenStudy (anonymous):

i think i understand it all :) thanks :)